Question 3 RawDeal is the new sushi bar in the neighborhood. Their estimated marginal cost is 10 cents per sushi unit. RawDea.
Answer #1

1) the marginal cost given is 10

the demand function given as Q= 20 - 10P so the total revenue will be

   TR = Q*P = (20 - 10P) * P = 20P - 10P2

   MR = d(TR)/ d (p) = 20 - 20P

now to get the equilibrium price per unit we need to equalise both the MC and MR

MC = MR

   10 = 20 - 20P

20P = 10

P = 10/20 = 0.5 cents per unit sushi.